OpenStudy (anonymous):

find the two numbers whose sum is 12, if the product of one by the square of the other is to be maximum a. 6 and 6 b. 5 and 7 c. 4 and 8 d. 9 and 3

OpenStudy (anonymous):

We'll note the numbers as x and y. The sum of the nubmbers is 12. x + y = 12 y = 12 - x We also know that the product of one and the square of the other is a maximum. We'll write the product of numbers as: P = x*y^2 We'll substitute y by (12-x) and we'll create the function p(x): p(x) = x*(12-x)^2 We'll expand the square: p(x) = x(144 - 24x + x^2) We'll remove the brackets: p(x) = 144x - 24x^2 + x^3 The function p(x) is a maximum when x is critical, that means that p'(x) = 0 We'll calculate the first derivative for p(x): p'(x) = (144x - 24x^2 + x^3)' p'(x) = 144 - 48x + 3x^2 p'(x) = 0 144 - 48x + 3x^2 = 0 We'll divide by 3 and we'll apply symmetric property: x^2 - 16x + 48 = 0 We'll apply the quadratic formula: x1 = [16+sqrt(256-192)]/2 x1 = (16+8)/2 x1 = 12 x2 = (16-8)/2 x2 = 4 The positive numbers are: x = 4 and y = 12 - 4 = 8

OpenStudy (anonymous):

which is c .

OpenStudy (anonymous):

wow u got it and nice explaination

OpenStudy (anonymous):

is it easier to substitute x? x + y = 12 xy^2 = p(x) (12 - y)y^2 = p(x) p'(x) = -y^2 + 2y(12 - y^2) -3y^2 + 24y = 0 y(24 - 3y) = 0 y = 8 x = 4

OpenStudy (anonymous):

w8 9 and 3 is also like that..

OpenStudy (anonymous):

3*81 = 243 4*64 = 256

OpenStudy (anonymous):

yes8 and 4

OpenStudy (anonymous):

ok he says 8 and 4

OpenStudy (anonymous):

u want the maxim, so 3*9^2=243 ,while 4*4^3=256

OpenStudy (anonymous):

plus we just proved that with maths

OpenStudy (anonymous):

the maximum in all letters is c =256 if i get the minimum the answer is the lower 1?

OpenStudy (anonymous):

i fyou said for minimum product, the answer would be 6, 6

OpenStudy (anonymous):

ty raza

OpenStudy (anonymous):

U WLCM